OpenStudy (anonymous):

Simplify, showing your steps. Tell the property you used in each step. 2(a + 4) + -8 i dont understand:(

OpenStudy (unklerhaukus):

distribute the two across the bracketed terms

OpenStudy (anonymous):

how?

OpenStudy (unklerhaukus):

\[2\times(a + 4)=2\times a+2\times4\]

OpenStudy (anonymous):

ohhh! okay i get that step now

OpenStudy (unklerhaukus):

that step was called distributing

OpenStudy (unklerhaukus):

can you finish form here?

OpenStudy (johnweldon1993):

2(a + 4) + -8 As UnkleRaukus has shown...you 'distribute' the 2 (multiply it by each term in the parenthesis) 2 * a = 2a + 2 * 4 = 8 so we have 2a + 8 - 8 = ...?